OpenStudy (anonymous):

@tHe_FiZiCx99 Which equation could be used to solve this problem? The sum of two consecutive integers is 141. Find the integers. A. n + 1= 141 B. 2n + 2 = 141 C. 2n = 141 D. 2n + 1 = 141

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@Rina.r

OpenStudy (agent0smith):

call the first n. The second is then n+1 (since they're consecutive). The sum is 141 n + n+1 = 141 simplify.

OpenStudy (anonymous):

so its A

OpenStudy (the_fizicx99):

2n + 2 = 141 do the rest

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

omg its A right

OpenStudy (the_fizicx99):

n + n+1 = x whatever x is in your case its 141

OpenStudy (anonymous):

its 4 question with 141!!

OpenStudy (agent0smith):

It is not A!

OpenStudy (agent0smith):

n + n+1 = 141 what is n+n?

OpenStudy (anonymous):

omg i dont know thats y i ask yall for help!!!!

OpenStudy (agent0smith):

what is 1n+1n?

OpenStudy (anonymous):

2

OpenStudy (agent0smith):

1n+1n is 2n

OpenStudy (anonymous):

yes

OpenStudy (agent0smith):

n + n+1 = 141 so... simplify. You know n+n is 2n.

OpenStudy (anonymous):

so its B

OpenStudy (agent0smith):

n+n+1 = 141 replace the n+n with 2n, what does the equation become?

OpenStudy (the_fizicx99):

x + (x + 1)

OpenStudy (anonymous):

2n + 2 = 141

OpenStudy (agent0smith):

Replace ONLY the n+n with 2n. Change NOTHING else. n+n +1 = 141

OpenStudy (anonymous):

omg so is it D

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

so who should get the medal

OpenStudy (the_fizicx99):

x+(x+1)=141 x+x+1=141 2x+1=141 -1 -1 2x=140 :)

OpenStudy (agent0smith):

@tHe_FiZiCx99 the question is not asking for that. See the options.

OpenStudy (anonymous):

omg omg omg omg

OpenStudy (the_fizicx99):

@agent0smith 2n + 1 = 141 Same, n is an unknown value it can be represented by any variable

OpenStudy (agent0smith):

@tHe_FiZiCx99 not my point. Look at your answer, 2x=140. Then look at the options.